### 14 Sep 2024 v2

# An Audit on the Perspectives and Challenges of Hallucinations in NLP

Abstract
We audit how hallucination in large language models (LLMs) is characterized in peer-reviewed literature, using a critical examination of 103 publications across NLP research. Through the examination of the literature, we identify a lack of agreement with the term ‘hallucination’ in the field of NLP. Additionally, to compliment our audit, we conduct a survey with 171 practitioners from the field of NLP and AI to capture varying perspectives on hallucination. Our analysis calls for the necessity of explicit definitions and frameworks outlining hallucination within NLP, highlighting potential challenges, and our survey inputs provide a thematic understanding of the influence and ramifications of hallucination in society.
1 Introduction
Recent advancements in Natural Language Processing (NLP) have expanded beyond traditional Machine Learning tools, evolving into sociotechnical systems that combine social and technical aspects to achieve specific goals Gautam et al. (2024); Narayanan Venkit (2023). They have now become integral in various domains such as health, policy-making, and entertainment, Jin and Mihalcea (2022); Werning (2024) showcasing their significant impact on daily life. However, language models (LM) exhibit negative behaviours such as hallucination and biases Bender et al. (2021); Gupta et al. (2024). This has catalyzed a surge in research investigating the phenomenon of hallucinations in NLP Ji et al. (2023a), reflected in the escalating number of peer-reviewed publications on the topic, as illustrated in Fig 1, sourced from SCOPUS.
Figure 1: Articles published each year (from 2013 to 2023) in SCOPUS that contain the term ‘hallucination’ AND (‘NLP’ OR ‘AI’) in the title, abstract, or keywords.
Within the NLP domain, various frameworks have emerged to define hallucination, primarily emphasizing its negative aspect. Hallucination here refers to the model’s production of references to non-existent objects or statements, lacking supporting examples in the training data Ji et al. (2023a). Despite the growing research on this topic, there is still a notable divide in our understanding, a lack of a unified framework, and a need for precise definitions Filippova (2020a).
The necessity to understand this gap is accentuated by research demonstrating the societal impacts of hallucinations Dahl et al. (2024). Hence, there is a growing need to explore how the field of NLP conceptualizes hallucination. In line with this imperative, the following questions guide this study:
  • RQ1: What are the definitions and frameworks used to explain hallucinations in NLP?
  • RQ2: What is the current understanding of researchers about hallucinations, and how do they encounter them in their work?
To answer the RQ1, we first conduct an audit of the field of hallucinations in NLP by surveying 103 peer-reviewed articles [^1] https://github.com/PranavNV/The-Thing-Called-Hallucination. Subsequently, we conduct a survey to 171 researchers and academics in the field to gather their perspectives on this phenomenon, providing a novel contribution to the literature, addressing RQ2. By surveying NLP practitioners, the paper incorporates real-world perspectives, enriching the theoretical discussions with practical insights. This audit therefore aims to broaden the communities perspective by presenting practical insights from researchers employing these methods in their work. We also propose an ethical framework to guide future efforts in comprehending and mitigating hallucinations in LLMs.
2 Evolution of Hallucination in NLP
The term ‘hallucination’ has a long history in machine learning and has been used in various contexts prior to the LM era. Its earliest documented usage can be traced to the 2000s when Baker and Kanade (2000) applied it in the context of image resolution enhancement, referring to the generation of new pixel values. Subsequently, "hallucination" has been frequently employed in computer vision research, including notable works such as Hsu et al. (2010) on face hallucination.
In the modern deep learning era, hallucination was used first by Andrej Karpathy in his blog focusing on Recurrent Neural Networks Karpathy (2015). He used the term within the context of LM by illustrating how an LSTM could generate non-existent URLs, effectively ‘hallucinating’ data. The term then gained major traction with the launch of ChatGPT Wu et al. (2023), where it referred to inaccuracies and factual mistakes produced by models Ji et al. (2023a). However, the field lacked a unified definition, leading to a spectrum of interpretations Filippova (2020b). In one of the earlier works, Maynez et al. (2020) divides term usage into intrinsic and extrinsic hallucination. Intrinsic hallucinations are consequences of synthesizing content using the information present in the input. Extrinsic hallucinations are model generations that ignore the source material altogether.
However, there is a rise in discussion around terminology that reflects a deeper inquiry into the phenomena, with recent discourse advocating for ‘confabulation’ Millidge (2023) or ‘fabrications’ McGowan et al. (2023) as a more precise descriptor. This reflects the lack of consensus on the term and highlights the importance of looking at the use of hallucination with a more critical lens.
3 Related Surveys on Hallucination
We now provide an overview of several key surveys in the realm of NLP focusing on the topic of hallucination and why our work addresses a relevant gap in the field. Starting with Ji et al. (2023a), this survey extensively delves into the advancements and challenges concerning hallucination in NLG, distinguishing between intrinsic and extrinsic frameworks of hallucination. Additionally, it sheds light on fundamental terms such as hallucination, faithfulness, and factuality, along with prevalent metrics for quantifying these phenomena. Rawte et al. (2023b) categorizes existing works within the domain of LMs, covering various aspects including methods for detecting hallucination, mitigation techniques, datasets used, and evaluation metrics. Zhang et al. (2023c) addresses the challenges of hallucination in LLMs by categorizing hallucinations into input-conflicting, context-conflicting, and fact-conflicting types, diverging from traditional viewpoints.
Furthermore, Huang et al. (2023) redefines the taxonomy of hallucination into factuality and faithfulness, with additional subdivisions, and proposes mitigation strategies aligned with underlying causes. Tonmoy et al. (2024) offer a comprehensive overview of over thirty-two techniques developed to mitigate hallucination in LLMs and finally, Rawte et al. (2023a) present a nuanced categorization of hallucination into six types, contributing to the ongoing discourse within the field.
While these surveys offer insights into the current state of hallucination research, they do not pay attention to critical examinations of the field’s weaknesses arising from a lack of discourse in defining hallucination and challenges due to the same. This deficiency in discussion reflects the broader trend within the entire field. Therefore, our audit answers this gap by critically examining how we conceptualize hallucination. We aim to highlight the challenges stemming from these definitions and to further conduct a practitioner survey within the community to understand researchers’ and developers’ perspectives on this issue.
4 Critical Analysis of Hallucination in NLP Literature
This section is dedicated to conducting an audit of hallucination research within NLP, aiming to uncover its applications and subsequently identify the strengths and weaknesses in current literature.
To accomplish this, we conducted an audit of works from the ACL anthology using specific keywords such as ‘hallucination’, ‘NLP (OR) AI’ AND ‘hallucinations’, ‘fabrication’, and ‘confabulations’. We surveyed papers released on and before April 19th, 2024. From this search, a total of 164 papers were retrieved. After filtering out papers that were not directly related to hallucination research or those that merely mentioned the term without substantial focus on the topic, we arrived at a corpus of 103 papers. This corpus forms the basis for our audit and analysis of hallucination research, specifically within the NLP domain.
4.1 Conceptualization of Hallucination
We performed an iterative thematic analysis Vaismoradi et al. (2013) to uncover the various applications of hallucination research in NLP. To ensure accuracy and prevent misclassification, this recursive process was employed. This resulted in the identification of seven distinct fields that address research on hallucination (as shown in Table. 1).
NLP Tasks Frequency
Conversational AI 38
Abstractive Summarization 16
Data-to-Text Generation 14
Machine Translation 12
Image-Video Captioning 8
Data Augmentation 8
Miscellaneous 7
Table 1: Frequency of papers reviewed for each thematically grouped NLP tasks.
This taxonomy affords insights into the pervasive nature of hallucination in NLP. Notably, it reveals that hallucination transcends beyond text generation, extending its conceptualization to encompass broader domains such as Image-Video Captioning, Data Augmentation, and Data-to-Text Generation tasks. This depicts the significance of hallucination both within and beyond the realm of NLP. Moreover, our classification framework provides us with a faceted analysis of how each of these tasks defines the concept of hallucination.
Using thematic categorization, we come across definite attributes across the definitions of hallucination. One set of attributes elucidated how hallucinations are associated with the style/language generated by the model: Fluency, Plausibility, and Confidence. The next set of attributes falls under the effects of hallucinations: Intrinsic, Extrinsic, Unfaithfulness and Nonsensical. The definition of each of these attributes is elaborated in Table 2.
In each paper analyzed within the survey scope, hallucination is defined based on a combination of the set of attributes identified. Our survey revealed 31 unique frameworks for conceptualizing hallucination, illustrating the diverse approaches and perspectives used. This diversity underscores the ambiguity in the term’s usage.
To illustrate this phenomenon, we present some examples showcasing the diverse approaches commonly observed in the literature:
[…]
Hallucination refers to the phenomenon where the model generates false information not supported by the input. - Xiao and Wang (2021a)
[…]
LLMs often exhibit a tendency to produce exceedingly confident, yet erroneous, assertions commonly referred to as hallucinations. - Zhang et al. (2023a)
[…]
Models generate plausible-sounding but unfaithful or nonsensical information called hallucinations - Ji et al. (2023c)
Hence, within NLP, a notable deficiency persists in grasping coherent characteristics of hallucination. This shortfall underscores the risk of potential misappropriation of the term when employed in divergent contexts. An extensive analysis of hallucination for each of the mentioned NLP tasks and its definition is illustrated in the Appendix 10.2.
Attributes Definition
Fluency
The syntactic incorrectness and semantic errors of the sentence generated.
Plausibility
The degree to which the generated text
appears factually incorrect or unbelievable within the given context.
Confidence
The absence of modifiers or qualifiers that express
uncertainty in the generated text, presenting the output with a sense of assuredness.
Intrinsic
The generated output that contradicts the source content or the input provided
Extrinsic
The generated output that cannot be verified from the source content or the input provided.
Non-factual
Inconsistent with facts in the real world, leading to the generation of non-factual
content in accordance to the established real-world knowledge.
Unfaithfulness
Inconsistent to the input prompt or context, creating deviations
or inconsistencies that would diverge from the intended meaning or message.
Nonsensical
Lack of logical meaning or coherence within a given context as well as the readability of the text.
Table 2: The attributes that appear in the definitions of hallucination.
4.2 Audit of Frameworks
We now scrutinize the dominant frameworks employed in defining hallucination while also assessing the extent to which these models accurately capture the phenomenon. We start by looking at how many of the selected works explicitly define hallucination. Out of the 103 papers, just 44 (42.7%) provide a definition of the term, leaving the majority—59 papers or 57.3%—either altogether omitting their understanding of hallucination in the context of their research or providing no definition or a framework. This lack of transparency is not only concerning but also underscores the need for clarity, especially given the varied interpretations of hallucination across different research domains.
Taking our scrutiny a step further, we investigate whether the works defining hallucination reference and acknowledge preexisting frameworks. It emerges that only 29 papers or 27% of the selected works explicitly acknowledge and adhere to established frameworks of hallucination, while the remainder 73% either loosely define the term or devise new definitions tailored to their specific research scope. This trend within the field shows a lack of consensus on the conceptualization of hallucination, leading to disparate interpretations and a shortage of discourse on the subject.
We also audit the sociotechnical nature of the definitions of hallucination in NLP. Hallucination (elucidated in Appendix 10.1) inherently contains social dimensions, creating varied perspectives across different social contexts. Moreover, given the evolution of LMs into social spaces, adopting a sociotechnical approach becomes necessary, given that the term ‘hallucination’ is inherently a shared vocabulary within these domains. Unfortunately, out of the 103 works examined, only 3 acknowledge the this nature of hallucination, with none utilizing this framework to inform their approach. This underscores a need for research to explore the sociotechnical dimensions inherent in hallucination, showcasing the limited depth of understanding within the ML and NLP communities.
4.3 Audit of Metrics
In the analysis of the 103 papers, we observed that 87 of these works dedicate efforts to measuring ‘hallucination.’ This observation depicts the prevailing trend within NLP, emphasizing the significance of quantifying the concept of hallucination across diverse research efforts. Building upon prior studies such as Ji et al. (2023a), our analysis categorizes the common approaches in NLP for quantifying hallucination into four major themes: Statistical Metrics, Data-driven Metrics, Human Evaluation, and Mixed Methodologies.
Figure 2: Hallucination evaluation metrics used in NLP.
Statistical metrics calculates a hallucination score based on the degree of mismatch, with higher discrepancies indicating lower accuracy, factuality or faithfulness and hence, higher hallucination Ji et al. (2023a). Statistical scores such as BLUE, ROUGE, and Error Rate metrics are commonly used in this approach. Our findings reveal that 35.2% of the works that quantify hallucination opt for statistical metrics, employing 25 distinct metrics (e.g., BERTScore, F1, Perplexity, Cosine Similarity) developed for this purpose. This variability underscores the lack of a standardized approach.
Data-driven metrics utilizes curated datasets or neural models to gauge hallucination in generated text. This methodology, accounting for curated knowledge/content mismatches, is adopted by 26.1% of the works, resulting in the development of 18 distinct datasets or models tailored for hallucination measurement, such as CHAIR (Caption Hallucination Assessment with Image Relevance) and SelfCheckGPT Manakul et al. (2023).
Human evaluation offers a complementary perspective by employing human annotators to assess hallucination levels, compensating for apparent errors in automated indicators Ji et al. (2023a). This approach, used by 10.2% of the works, encompasses scoring and comparison methods, where annotators rate hallucination levels or compare output texts with baselines or ground-truth references. Notably, one outlier paper introduced an innovative approach utilizing eye tracking for hallucination detection in NLP tasks Maharaj et al. (2023).
Mixed method approach is deployed by 28.4% of the works, combining human evaluation with statistical metrics to offer a holistic perspective on hallucination quantification. This trend reflects a concerted effort within the research community to address the limitations of individual methodologies and provide insights into the presence and nature of hallucination in generated texts.
The metrics audit reveals significant knowledge gaps and challenges across various approaches. Notably, established research highlights areas for improvement in standard methods for measuring hallucination. For instance, methodologies like CHAIR and metrics such as ROUGE scores exhibit instability in measuring hallucination due to the need for complex human-crafted parsing rules for exact matching, rendering them susceptible to errors Li et al. (2023). Criticisms also extend to human evaluation methods, which are prone to inaccuracies in gauging hallucination within these models Smith et al. (2022).
Beyond methodological criticisms, our audit uncovers a trend of employing numerous distinct metrics and approaches within these frameworks to categorize hallucinations. Over time, this has led to a diverse set of parameters for measuring hallucination, with a general lack of consensus on a standardized measurement approach. This issue further highlights the absence of a unified method, especially as these models have now shifted to become a sociotechnical solution Bender et al. (2021).
5 Practitioner Survey of Hallucination
In this section, adopting a ‘community-centric approach’ Narayanan Venkit (2023), we conduct a survey to gain insights into researchers’ perceptions of hallucinations in NLP to complement our theoretical discussions with practical real-world perspectives. The primary goal is to demonstrate how researchers and practitioners within the field perceive the concept of ‘hallucination’ and to expand our findings beyond the limitations of existing literature where real-world perceptions from the researchers are missing Huang et al. (2023); Zhang et al. (2023c); Ji et al. (2023a). This motivates us to gather real-world perspectives from individuals actively engaged in NLP and AI research.
5.1 Survey Recruitment and Data Collection
For our survey, we employed a multi-faceted approach to reach a diverse population of respondents. We utilized direct emails, direct messages, and social media platforms such as LinkedIn and Twitter to distribute the survey. Our target audience included graduate students and professors from academic backgrounds as well as individuals from the industry who work in NLP, aiming to capture a wide range of perspectives on hallucinations.
To ensure a comprehensive view, we specifically targeted researchers familiar with AI and ML, primarily from disciplines such as computer science and information science. However, we also welcomed participants from other domains to explore their perceptions of whether they had the literature understanding of the concept of hallucination as they are also extensively using LLM models. The survey was examined and approved by the Institutional Review Board (IRB) for ethical practices.
We additionally employed a systematic approach by randomly selecting 15 universities from the top 100 in the USA as per the 2023 US News and World Report rankings News (2023), to then reach out to potential participants. Prior works Chakravorti et al. (2023) have previously employed this process to identify high-quality participants. We received a total of 223 responses, out of which 171 were complete and usable for analysis.
5.2 Survey Structure
The survey employed a combination of 14 open-ended and close-ended questions. The survey has been built based on the previous survey design techniques Rosen et al. (2013); Baker (2016); Van Noorden and Perkel (2023); Chakravorti et al. (2024). Open-ended questions and free-response text boxes allow us to gather rich opinions from participants. This approach integrates all our findings, providing a broader and deeper understanding of the response. For the analysis of open-ended questions, we utilized thematic analysis, drawing from established methodologies outlined in Blandford et al. (2016); Terry et al. (2017). The close-ended questions were analyzed using descriptive statistics to summarize and analyze the numerical data obtained from respondents. Throughout the analysis process, the research team made collective decisions regarding the retention, removal, or reorganization of themes derived from open-ended responses. All the survey questions have been provided in Appendix 10.4.
5.3 Survey Findings
We now summarize insights from our responses to explore various perspectives on hallucinations in LLMs, including perceptions, weaknesses, and preferences. The breakdown of responses indicates that 76.54% of participants were from academia, 20.98% from the industry, and 2.47% both.
Participants were also asked about their research area’s direct relation to AI and NLP. The analysis revealed that more than 68.52% of researchers indicated that their work is directly related to NLP, while the remaining respondents either exhibited familiarity with or indirectly incorporated NLP and AI methodologies in their work. This highlights the substantial involvement of AI experts and practitioners within the survey.
5.3.1 Familiarity with Hallucination
The survey included the question on participants’ familiarity with the concept of ‘hallucinations’ in AI-generated text, measured on a 5-point Likert scale. The analysis revealed that 24.07% of researchers reported being extremely familiar with the concept, while 33.33% indicated being very familiar with it (Figure 3). Participants who indicated not being familiar with the term ‘hallucination’ (7.96%) also demonstrated implicit concerns with this phenomenon by highlighting issues such as generating incorrect responses and crafting stories autonomously. This demonstrates the widespread impact of the phenomenon within the community.
Figure 3: Respondents familiarity with ‘Hallucination’
Figure 4: Frequency of encountering ‘Hallucination’
5.3.2 Hallucination Frequency
The survey included a question regarding the frequency of encountering ‘hallucinated’ content, defined as content that is factually incorrect or unrelated to the input, assessed on a 5-point Likert scale ranging from ‘Never’ to ‘Very frequently’ (Figure 4). The analysis revealed that 46.91% of respondents reported encountering hallucinated content occasionally, while 29.01% indicated experiencing it frequently. The results suggest that a substantial portion of practitioners encounter instances of hallucinated content in AI-generated outputs, indicating a prevalent issue in generative NLP models.
5.3.3 Perceptions of Hallucination
The survey findings revealed that more than 92% of respondents perceive hallucination as a weakness of LLMs. Subsequently, participants were asked to provide their own definitions of ‘hallucination’ in generative AI models through an open-ended question. To analyze these responses, we applied thematic categorization based on attributes generated from the literature audit (Table 2).
The thematic categorization revealed that the majority of respondents categorized hallucination as pertaining to the factuality and faithfulness of input, with relatively lesser emphasis on the extrinsic and intrinsic nature of hallucination concerning the input. This trend reflects a common perception of how hallucination is understood within the context of larger-scope generative AI models.
Moreover, the analysis identified 12 distinct frameworks regarding how hallucination is defined by respondents. For example:
[…]
Response that appears syntactically and semantically believable, but is not based on actual fact—Academic Researcher, NLP
[…]
When the model confidently states something that is not true—Academic Researcher, AI
The diversity of viewpoints underscores the inconsistency within the field regarding the conceptualization and understanding of hallucination in the context of generative AI models.
5.3.4 Alternative Terms for Hallucination
The survey included a question asking participants if they prefer an alternate term to describe the phenomenon of ‘hallucination’ in AI-generated content and to provide an explanation if they do. The analysis revealed that 54.32% of respondents preferred the term hallucination or had no other term to provide. However, among the remaining responses, 40.46% of participants mentioned ‘Fabrication’ as a better term to describe the phenomenon.
This indicates that while the majority of respondents did not propose an alternative term, a notable proportion sees fabrications as a more suitable descriptor for the phenomenon of hallucination in AI-generated content. For example,
[…]
Fabrication makes more sense. Hallucination makes it feel like AI is human and has the same sensory perceptions that could lead to hallucinations.—Academic Researcher, AI & Education
It’s interesting to note that a few researchers also prefer to use the term ‘Confabulations’ instead of ‘hallucinations’ when referring to AI-generated content. Their rationale likely stems from the nuanced difference in meaning between the two terms. While hallucinations generally convey the idea of perceiving something that is not based on reality or fact, confabulations specifically refer to the creation of false memories or information without the intention to deceive.
By opting for the term ‘Confabulations,’ researchers may be emphasizing the unintentional nature of the inaccuracies or false information generated by AI models, as opposed to implying deliberate deceit. For example,
[…]
I think confabulation works better because it means creating a false memory without deceit. Fabrication gives the idea that it is intentional, which in the case of generative AI models, it is not.—Academic Researcher, AI & HCI
It’s also insightful to see that respondents proposed various alternative terms to describe the phenomenon of hallucination in AI-generated content such as incorrect information/misinformation, Non-factual information, Cognitive gap, hyper-generalization, Overconfidence, and Randomness. These alternatives highlight different aspects and nuances of the inaccuracies or distortions present in the generated content. Participants also mentioned how they prefer multiple terms based on the application in which they are used.
[…]
As I mentioned there are different types of hallucinations. For instruction and context hallucinations, I would refer to them as inconsistency instead. For factually incorrect hallucinations, the word hallucination is fine.—Academia, NLP
5.3.5 Creativity and Positive Applications
[…]
Hallucinations are just what is needed for models to be creative. In truth, unless AI text-generators are factually grounded with external knowledge for a specific field, they are just story generators which aim to be creative, hence“hallucinate."—National Lab Researcher, NLP
Further supplementary analysis and quotes on the various external perspectives and the societal ramifications of hallucination, obtained through the survey, is examined in the Appendix 10.3.
6 Challenges and Recommendations
Based on our audit and survey analysis, we outline the key weaknesses encountered in hallucination within NLP and potential recommendations motivated by the weaknesses. We utilize a community-centric approach to define the primary weaknesses of the field currently and a path forward.
6.1 Challenges
The primary challenges we identify thematically and aim to elucidate are as follows:
Wide range of vague and absent definitions: The literature and the practitioner’s survey show diverse and conflicting frameworks, often lacking clarity or omitting explicit definitions for hallucination and how it is perceived in various fields of NLP and language generation. Ambiguity arises from the use of terms like ‘confabulations,’ ‘ fabrications,’ ‘misinformation,’ and ‘hallucinations’ interchangeably, without clear definitions in the context of hallucinations.
Lack of standardization in measurement: The absence of standardized metrics to quantify hallucination results in the use of multiple scales and categorizations. This makes it challenging to compare and interpret results across different models and studies, leading to a proliferation of diverse approaches for evaluating hallucinations.
Limited awareness of hallucination in a sociotechnical context: Hallucination research often lacks the understanding of how the concept of hallucination is conceptualized beyond its technical purview. When such analysis is employed in sociotechnical systems like healthcare and policy making Dahl et al. (2024); Pal et al. (2023), it is important to define the socially relevant framework of hallucination as well. There is no motivation shown to understand the non-technical considerations of hallucination.
Multiple sentiment towards hallucination: The perception of hallucination in generative AI varies depending on the context. For instance, it is often positively regarded as creativity in image generation, whereas in text generation, it is viewed negatively as errors or mistakes. Consequently, future research efforts should aim to better address this disparity to develop a more nuanced framework for understanding hallucination.
Lack of standardized nomenclature: Both our literature audit and practitioner survey revealed that the term ‘hallucination’ is inadequate to fully capture the behavior exhibited by NLG models. There is a need for further investigation into which terms are more appropriate and why they are necessary. For instance, terms like ‘confabulation,’ ‘fabrications,’ and ‘misinformations’ are increasingly being used to describe the same phenomenon. A more precise understanding is required to distinguish between these terms and how they are utilized in various fields within NLP.
User trust and reliability: Our survey findings suggest that users may hesitate to fully utilize LLM capabilities due to concerns about bias and hallucination despite recognizing the potential advantages these models offer. Therefore, there is a need to focus efforts on understanding the human interaction aspect concerning hallucination in NLP and language generation.
Addressing these issues requires careful consideration of the categorization approach, integration of contextual information, and, efforts towards robust evaluation methodologies in hallucinations.
6.2 Recommendations
Expanding on audits like Blodgett et al. (2020) & Venkit et al. (2023), we examine strategies for NLP practitioners studying ‘hallucination’ to overcome these challenges. We propose two overarching themes with four associated recommendations.
Author-Centric Recommendation.
These recommendations prioritize actionable steps for both the author and developers, emphasizing transparent and accountable development in conceptualizing hallucinations.
[R1] Ensure explicit documentation of the hallucination framework and analysis methodology employed during the development of NLP models. Provide guidelines that outline the expected measurements and quantifications for the model to enhance interpretability and applicability.
[R2] Explicitly state the use cases and user profiles intended to interact with the NLP system. By considering the specific applications and targeted users, it is easier to construct the required framework of hallucination that is sensitive to the community in consideration. Raise awareness about potential ramifications introduced by NLP models, emphasizing the importance of fairness and equity.
Community-Centric Recommendation.
These recommendations prioritize actionable steps for the research community to enhance frameworks and understanding related to hallucinations.
[R3] Develop clear and standardized definitions for terms such as ‘confabulations,’ ‘fabrications,’ ‘misinformation,’ and ‘hallucinations’ within the context of NLP. Establish frameworks that provide clarity and consistency in understanding these concepts, particularly regarding hallucinations. This requirement is crucial due to the widespread misunderstanding of hallucination and the misnomers that have arisen as research progresses.
[R4] Promote the creation of methods that offer visibility into the model’s decision-making process, enabling users to comprehend how hallucinations or fabrications can occur within the system, thus fostering trust in its use. Facilitating research discussions for transparency through workshops and conferences is one approach to achieving this goal.
7 Conclusion
Our work delves into the conceptualization of hallucination within the scope of NLP. Our approach involved two key methodologies: first, an exhaustive audit of 103 peer-reviewed papers in the NLP domain, and second, a practitioner survey of 171 researchers to complement our first study with real-world practical perception and understanding of hallucination as a unique contribution. Through this analysis, we have gained insights into how the NLP community conceptualizes and defines hallucination, showcasing a lack of discourse and agreement. Additionally, our thematic and community-based approach highlights potential weaknesses within the field, particularly in addressing misrepresentations and inaccurate characterizations associated with hallucination, paving way for better advancement in language generation.
8 Limitations
Our study encompasses a selection of 103 papers, incorporating works from primarily the ACL Anthology. While our intention was not to provide an exhaustive collection of all published works on hallucination, we aimed to include diverse sources within NLP that cover various aspects of the field. Our intent was to curate peer-reviewed literature commonly found in the NLP domain, encompassing models, applications, survey papers, and frameworks. We, therefore, did not scope the utility of hallucination and its impact beyond NLP to other fields of research, such as Computer Vision. Regarding the creation of the challenges and recommendations, it is important to note that the themes presented are not meant to be exhaustive but rather serve as a foundational framework to spark additional inquiries and foster further engagement.
Our survey was designed to capture the viewpoints of researchers and practitioners in the AI and ML field, potentially limiting various experiences. As such, our analysis is centered on this perspective. While we did gather additional insights from participants outside this field, our focus was not comprehensive in that regard. Our future work intends to explore the public’s perspective on hallucination.
9 Ethics Statement
We are aware of the ethical considerations involved in our research and have taken measures to ensure responsible practices throughout the study.
Data Publication: All the papers used in our research are listed in the Appendix. However, we recognize the importance of transparency and accountability. Therefore, we publish the complete collection of papers along with our qualitative classification and annotation, allowing for public scrutiny and examination  [^2] https://github.com/PranavNV/The-Thing-Called-Hallucination.
Mitigating Qualitative Bias: We acknowledge the potential for bias when performing qualitative coding of the papers regarding their applications. To address this concern, we ensured that at least three different individuals independently reviewed and verified the coding to minimize the possibility of misclassification. Additionally, we followed the same approach to verify the presence of various definitions in each paper, enhancing the reliability and validity of our analysis. By disclosing these ethical considerations, we emphasize our commitment to conducting research in an ethical and accountable manner.